OpenStudy (cutiecomittee123):

In the following triangle, Theta=60. Find the value of the angle B1 and B2 which solve this ambiguous case

OpenStudy (cutiecomittee123):



OpenStudy (cutiecomittee123):

I know that we have to use law of sines So sin60/12.3=sinB/13 but, I tried solving this way and I came up with .9 something which doesn't satisfy the angle.

OpenStudy (dan815):

use degrees mode on your calculator

OpenStudy (cutiecomittee123):

It is in degrees

OpenStudy (cutiecomittee123):

Im getting .9153114024 and if I take sin inverse I get B=66.93 Oh I think I got it so then 180-66.93=113.07

OpenStudy (cutiecomittee123):

so b1=66.93 and B2=113.07